Three times Sally's age now exceeds 5 times John's age by 15. In 10 years, twice Sally's age then will be 20 less than 4 times John's age then. How old will they be 15 years from now?
The equations I have are this:
1st sentence: 3S = 5J +15
2nd sentence: 2(S + 10) = 4(J + 10) + 20

Answer by ankor@dixie-net.com(22740)

Three times Sally's age now exceeds 5 times John's age by 15. In 10 years, twice Sally's age then will be 20 less than 4 times John's age then. How old will they be 15 years from now?
The equations I have are this:
:
1st sentence: 3S = 5J + 15
this is right
:
2nd sentence: 2(S + 10) = 4(J + 10) + 20
this equation should be
2(S + 10) = 4(J + 10) - 20
2s + 20 = 4j + 40 - 20
2s + 20 = 4j + 20
2s = 4j + 20 - 20
2s = 4j
simplify, divide by 2
s = 2j
now substitute 2j for s in the 1st equation